How do I solve by addition method
3x-5y=5
x+2y=9

Answer by stanbon(75887)

How do I solve by addition method
3x-5y=5
x+2y=9
---------
Multiply thru the 2nd equation by 3:
3x+6y = 27
---
Subtract 3x-5y=5 from 3x+6y=27 to get:
11y = 22
y = 2
------
Substitute that into x+2y=9 to solve for "x":
x + 2*2 = 9
x = 5
--------------------
Cheers,
Stan H.